# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix C Relationship to Group DP-SGD
The dependence on the discrepancy cap $K_{t}$ (worst-case) and its mean-square analogue $\mathbb{E}_{\eta}[K_{t}^{2}]$ in our SRPP/ms-SRPP privatization of SGD naturally raises the question: can group DP-SGD, with an appropriate group size, serve as a reliable proxy for SRPP-SGD when the protected information is a specific Pufferfish secret? In this appendix, we clarify the relationship between SRPP-/ms-SRPP-SGD and group DP-SGD, and explain why the latter cannot generally substitute for the former when one cares about secret-level Pufferfish guarantees.
@@
Our SRPP- and ms-SRPP-SGD mechanisms are calibrated directly from the HUC and ms-HUC envelopes developed in Section LABEL:sec:srpp_sgd. For a fixed Rényi order $\alpha>1$ and a target sliced privacy budget $\epsilon$, the corresponding theorems give an upper bound on the per-run SRPP or ms-SRPP cost in terms of the HUC/ms-HUC sequence $\{h_{t}\}$ and the per-step noise covariances $\{\Sigma_{t}\}$. In our experiments we specialize to isotropic Gaussian noise $\Sigma_{t}=\sigma^{2}I$ and, for each desired $\epsilon$, solve the resulting bound for a noise scale $\sigma(\epsilon)$ that guarantees the prescribed SRPP or ms-SRPP budget. This calibration is explicitly secret-aware: $H_{\mathrm{total}}$ depends on the Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$ through the discrepancy caps $K_{t}$ and mean-square discrepancy caps $\mathbb{E}_{\eta}[K_{t}^{2}]$, which, for each iteration $t$, bound the $\ell_{2}$–distance between the clipped, averaged gradient updates under any admissible secret pair $(s_{i},s_{j})\in\mathcal{Q}$ and any coupling of their corresponding data-generating distributions.
@@
$\Pr[\mathcal{M}(X)\in E]\leq e^{\varepsilon_{g}}\Pr[M(X^{\prime})\in E]+\delta_{g},\quad\forall E\subseteq\mathcal{Y},\ \forall X\sim_{k}X^{\prime}.$ (25)
In DP-SGD, one typically fixes $k$ and uses a moments accountant to obtain $(\varepsilon_{g},\delta_{g})$ as a function of $k$, the sampling probability $q=L/N$, the number of steps $T$, and the per-step noise scale $\sigma_{\mathrm{DP}}$.
A natural question is whether group DP-SGD can be used as a proxy for SRPP-/ms-SRPP-SGD by setting the group size to $k=\mathbb{E}_{\eta}[K_{t}^{2}]$. In what follows, we focus on this mean-square quantity because using the worst-case cap $K_{t}$ as an effective group size would only further increase the required noise and make group DP even more conservative.
@@
  • There is no general result that, given a group-DP guarantee (25) for some group size $k$, produces an $(\alpha,\varepsilon)$-SRPP or ms-SRPP guarantee for an arbitrary Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$. In particular, even if we set $k=\mathbb{E}_{\eta}[K_{t}^{2}]$ based on our HUC/ms-HUC analysis, there is no closed-form or universal mapping from $(\epsilon_{g},\delta_{g},k)$ to an SRPP/ms-SRPP budget $\epsilon$.
  • Even when we fix a specific Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$ and choose the group size $k=\mathbb{E}_{\eta}[K_{t}^{2}]$, the resulting $(\epsilon_{g},\delta_{g})$ group-DP-SGD guarantee is in general not a certified conservative bound for our $(\alpha,\varepsilon)$ SRPP/ms-SRPP guarantees, and there is no theorem ensuring that this group-DP guarantee upper-bounds the leakage measured in the SRPP/ms-SRPP sense.
C.1 No Universal Mapping
We first formalize the absence of a universal conversion from group-DP guarantees to SRPP/ms-SRPP guarantees.
@@ A DP-to-SRPP conversion map is a function
that, for given group-DP parameters $(\epsilon_{g},\delta_{g})$ and group size $k$, assigns an SRPP budget $F(\epsilon_{g},\delta_{g},k)$ with the following property: for every mechanism $\mathcal{M}$ and every Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$, if $\mathcal{M}$ is $(\epsilon_{g},\delta_{g})$-group-DP with group size $k$, then for all $\theta\in\Theta$ and all $(s_{i},s_{j})\in\mathcal{Q}$ the SRPP/ms-SRPP divergence between the laws of $\mathcal{M}$ under $(S=s_{i},\theta)$ and $(S=s_{j},\theta)$ is bounded by $F(\epsilon_{g},\delta_{g},k)$.
@@
Proposition C.1.
There is no function $F$ satisfying Definition C.1, even if we restrict to pure group-DP ($\delta_{g}=0$) and to finite secret sets $\mathcal{S}$.
Proof of Proposition C.1.
@@
Let the dataset space be $\mathcal{X}=\{X_{0},X_{1}\}$, and choose $k\in\mathbb{N}$ so that $d_{H}(X_{0},X_{1})>k$, where $d_{H}$ is the Hamming distance. Under the group adjacency relation $X\sim_{k}X^{\prime}$ (defined by $d_{H}(X,X^{\prime})\leq k$), there are no distinct adjacent pairs: the only pairs with $d_{H}(X,X^{\prime})\leq k$ are $(X_{0},X_{0})$ and $(X_{1},X_{1})$. Hence, for any mechanism $\mathcal{M}$ and any $(\epsilon_{g},\delta_{g})$, the group-DP inequality (25) is trivially satisfied, so $\mathcal{M}$ is $(\epsilon_{g},\delta_{g})$-group-DP for all choices of $(\epsilon_{g},\delta_{g})$ and $k$.
@@
In particular, the Rényi divergence of any order $\alpha>1$ between the laws of $\mathcal{M}(X)$ given $S=0$ and $S=1$ is $\infty$, so no finite SRPP budget can hold.
@@
Proposition C.1 remains true if, instead of allowing arbitrary mechanisms, we restrict attention to mechanisms obtained by running DP-SGD and calibrating noise for a fixed group size $k$ (including the specific choices $k=K_{\mathrm{cap}}$ or $k\approx\mathbb{E}_{\eta}[K_{t}^{2}]$ from our HUC/ms-HUC analysis). We can always construct a Pufferfish scenario whose secrets are encoded in dataset differences at Hamming distance strictly greater than $k$ and hence outside the scope of the group-DP guarantee. Therefore, even when the group-DP parameters $(\epsilon_{g},\delta_{g},k)$ are known from a moments-accountant analysis of DP-SGD, there is no closed-form or universal transformation that converts them into the SRPP/ms-SRPP budget used in our HUC/ms-HUC–based SRPP-SGD calibration.
C.2 No Safe Bounds
We now specialize to the Pufferfish scenarios used in our experiments of SGD privatization, where the secret encodes the prevalence of a protected class (e.g. the cat label in CIFAR-10). Concretely, we consider two secret values $s_{0},s_{1}$ which correspond to two label configurations on the same feature set, with different prevalences $p_{\mathrm{low}}$ and $p_{\mathrm{high}}$ of the protected class. Let $X^{(0)}$ and $X^{(1)}$ denote the realized datasets under $S=s_{0}$ and $S=s_{1}$, respectively, and let
@@ C.2 No Safe Bounds
be the number of points whose labels differ between the two secrets.
In our HUC/sm-HUC analysis, $\tilde{\Delta}$ enters explicitly through the random variable $K_{t}$ (the number of “differing” examples in a mini-batch at step $t$) and its cap $K_{\mathrm{cap}}$ (the worst-case throughout) or second moment $\mathbb{E}_{\eta}[K_{t}^{2}]$, which yield a scalar $\hat{\gamma}^{2}$ and hence $H_{\mathrm{total}}$ for SRPP/ms-SRPP accounting. One might ask whether it is sufficient, instead, to treat such quantities as an effective group size and calibrate a standard group DP-SGD with $k=\mathbb{E}_{\eta}[K_{t}^{2}]$. The following proposition shows that, even in prevalence-based scenarios, such a group-DP guarantee does not induce a general conservative (i.e., “safe") bound for the secret.
Proposition C.2.
Fix any group size $k\in\mathbb{N}$ and any group-DP parameters $(\varepsilon_{g},\delta_{g})$. There exist:
@@
    a mechanism $\mathcal{M}$ that is $(\epsilon_{g},\delta_{g})$-group-DP with group size $k$
such that the SRPP/ms-SRPP leakage between $(S=s_{0},\theta)$ and $(S=s_{1},\theta)$ under $\mathcal{M}$ is infinite (in particular, not bounded by any finite function of $(\epsilon_{g},\delta_{g},k)$).
@@
Fix $k\in\mathbb{N}$ and $(\epsilon_{g},\delta_{g})$. Choose a dataset size $N>k$ and an integer $\tilde{\Delta}\in\{k+1,\dots,N\}$. We construct a prevalence-based Pufferfish scenario and a mechanism $M$.
@@
Let the data domain be $\bar{\mathcal{X}}=\mathcal{X}\times\{0,1\}$, where the second coordinate is a binary label indicating membership in the protected class (e.g. cat vs. “not-cat”). Let $\mathcal{X}$ be the space of datasets of size $N$ over $\bar{\mathcal{X}}$: $\mathcal{X}:=\bar{\mathcal{X}}^{N}$. Fix two label configurations (datasets) $X^{(0)},X^{(1)}\in\mathcal{X}$ on the same feature vectors such that:
  • (i)
    the empirical prevalence of the protected class in $X^{(0)}$ is $p_{\mathrm{low}}$, and in $X^{(1)}$ is $p_{\mathrm{high}}$, with $p_{\mathrm{high}}\neq p_{\mathrm{low}}$, and
  • (ii)
    exactly $\Delta$ labels differ between $X^{(0)}$ and $X^{(1)}$, so that
    $d_{H}(X^{(0)},X^{(1)})=\Delta>k$
For example, we may take $X^{(0)}$ to have $\lfloor p_{\mathrm{low}}N\rfloor$ protected labels and $X^{(1)}$ to have $\lfloor p_{\mathrm{high}}N\rfloor$ protected labels, obtained by flipping exactly $\Delta$ labels.
Define a Pufferfish scenario with:
$\mathcal{S}=\{s_{0},s_{1}\},\quad\Theta=\{\theta\},\qquad\mathcal{Q}=\{(s_{0},s_{1}),(s_{1},s_{0})\}.$
The data-generating mechanism is deterministic:
$P_{\theta}(X=X^{(0)}\mid S=s_{0})=1,\quad P_{\theta}(X=X^{(1)}\mid S=s_{1})=1.$
Thus, the secret $S$ selects which prevalence configuration (dataset) is realized.
Group-DP adjacency and mechanism.
Consider the group-DP adjacency relation on $\mathcal{X}$ defined by Hamming distance at most $k$:
@@
Since $d_{H}(X^{(0)},X^{(1)})=\Delta>k$, the pair $(X^{(0)},X^{(1)})$ is not adjacent. In particular, the group-DP constraint (25) never compares the joint law of $(\mathcal{M}(X^{(0)}),\mathcal{M}(X^{(1)}))$.
We now construct a mechanism $\mathcal{M}$. First, pick any mechanism $\mathcal{M}_{0}:\mathcal{X}\to\mathcal{Y}$ that is $(\epsilon_{g},\delta_{g})$-group-DP with group size $k$ (for example, the constant mechanism $\mathcal{M}_{0}(X)\equiv y_{*}$ for some $y_{*}\in\mathcal{Y}$). Such a mechanism exists for all $(\epsilon_{g},\delta_{g},k)$.
Define a new mechanism $\mathcal{M}$ by modifying $\mathcal{M}_{0}$ only on the two special datasets $X^{(0)}$ and $X^{(1)}$:
$\mathcal{M}(X):=\begin{cases}y_{0},&\text{if }X=X^{(0)},\\ y_{1},&\text{if }X=X^{(1)},\\ M_{0}(X),&\text{otherwise},\end{cases}$
where $y_{0},y_{1}\in\mathcal{Y}$ are two distinct outputs such that $\mathcal{M}(X^{(0)})$ and $\mathcal{M}(X^{(1)})$ are almost surely separated (for example, take deterministic outputs $y_{0}\neq y_{1}$).
We claim that $M$ is still $(\epsilon_{g},\delta_{g})$-group-DP with group size $k$. To see this, note that the definition of group-DP requires (25) to hold for all adjacent pairs $X\sim_{k}X^{\prime}$, i.e. all pairs with $d_{H}(X,X^{\prime})\leq k$. We distinguish two cases:
  • (i)
    If $X\sim_{k}X^{\prime}$ and $\{X,X^{\prime}\}\cap\{X^{(0)},X^{(1)}\}=\emptyset$, then $\mathcal{M}(X)=\mathcal{M}_{0}(X)$ and $\mathcal{M}(X^{\prime})=\mathcal{M}_{0}(X^{\prime})$, so the inequality (25) holds for $\mathcal{M}$ because it holds for $\mathcal{M}_{0}$.
  • (ii)
    If $X\sim_{k}X^{\prime}$ and, say, $X=X^{(0)}$, then $d_{H}(X^{(0)},X^{\prime})\leq k$ and $X^{\prime}\neq X^{(1)}$ (since $d_{H}(X^{(0)},X^{(1)})=\Delta>k$). By construction, $\mathcal{M}$ agrees with $\mathcal{M}_{0}$ on $X^{\prime}$, and we can choose the distribution of $\mathcal{M}(X^{(0)})$ so that (25) holds for all such adjacent pairs; for example, we can set $\mathcal{M}(X^{(0)})\stackrel{{\scriptstyle d}}{{=}}\mathcal{M}_{0}(X^{(0)})$ and then only encode the secret in the choice of labels $(X^{(0)},X^{(1)})$ in the Pufferfish scenario, rather than in the law of $\mathcal{M}(X^{(0)})$ itself. An entirely analogous argument applies if $X=X^{(1)}$.
Thus $\mathcal{M}$ can be constructed so that it remains $(\epsilon_{g},\delta_{g})$-group-DP with group size $k$ and yet its behavior on the two realized datasets $X^{(0)}$ and $X^{(1)}$ is separated as much as we like.
SRPP/ms-SRPP leakage.
Under the Pufferfish scenario above, the distribution of $\mathcal{M}(X)$ given $S=s_{0}$ is the law of $\mathcal{M}(X^{(0)})$, and the distribution of $\mathcal{M}(X)$ given $S=s_{1}$ is the law of $\mathcal{M}(X^{(1)})$. By choosing $\mathcal{M}$ so that these two laws are mutually singular (e.g. deterministic, distinct outputs), we obtain
$\Pr[\mathcal{M}(X)\in E\mid S=s_{0}]=1,\qquad\Pr[\mathcal{M}(X)\in E\mid S=s_{1}]=0$
for some measurable set $E\subseteq\mathcal{Y}$. Hence the Rényi divergence of any order $\alpha>1$ between the laws of $\mathcal{M}(X)$ under $S=s_{0}$ and $S=s_{1}$ is $\infty$, so the SRPP (and ms-SRPP) leakage between these two secrets is infinite. In particular, it cannot be bounded by any finite function of $(\varepsilon_{g},\delta_{g},k)$.
Since $(\varepsilon_{g},\delta_{g},k)$ were arbitrary, this establishes the claim. ∎
@@
Proposition C.2 shows that, even when we work with prevalence-based Pufferfish secrets as in our CIFAR-10 experiments, simply calibrating DP-SGD with a group size chosen from our HUC/ms-HUC analysis (for example $k=K_{\mathrm{cap}}$ or $k\approx\mathbb{E}_{\eta}[K_{t}^{2}]$) does not yield a certified bound in the SRPP/ms-SRPP sense. Group-DP is a guarantee indexed by Hamming-adjacent dataset pairs, whereas SRPP/ms-SRPP explicitly quantify the divergence between output distributions induced by different secrets (and different priors) in a given Pufferfish scenario. Our HUC/sm-HUC– based SRPP-SGD calibration is therefore not a reparameterization of group DP-SGD; it is a distinct analysis that targets secret-level leakage for the specified Pufferfish scenario.
Appendix D Minimality and Attainability of HUC
Proposition D.1.

### v3

@@ Appendix G Relationship to Group DP-SGD
The dependence on the discrepancy cap $K_{t}$ (worst-case) and its mean-square analogue $\mathbb{E}_{\eta}[K_{t}^{2}]$ in our SRPP/sa-SRPP privatization of SGD naturally raises the question: can group DP-SGD, with an appropriate group size, serve as a reliable proxy for SRPP-SGD when the protected information is a specific Pufferfish secret? In this appendix, we clarify the relationship between SRPP-/sa-SRPP-SGD and group DP-SGD, and explain why the latter cannot generally substitute for the former when one cares about secret-level Pufferfish guarantees.
@@
Our SRPP- and sa-SRPP-SGD mechanisms are calibrated directly from the HUC and sa-HUC envelopes developed in Sections 5.1 and 5.2, respectively. For a fixed Rényi order $\alpha>1$ and a target sliced privacy budget $\epsilon$, the corresponding theorems give an upper bound on the per-run SRPP or sa-SRPP cost in terms of the HUC/sa-HUC sequence $\{h_{t}\}$ and the per-step noise covariances $\{\Sigma_{t}\}$. In our experiments we specialize to isotropic Gaussian noise $\Sigma_{t}=\sigma^{2}I$ and, for each desired $\epsilon$, solve the resulting bound for a noise scale $\sigma(\epsilon)$ that guarantees the prescribed SRPP or sa-SRPP budget. This calibration is explicitly secret-aware: $H_{\mathrm{total}}$ depends on the Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$ through the discrepancy caps $K_{t}$ and mean-square discrepancy caps $\mathbb{E}_{\eta}[K_{t}^{2}]$, which, for each iteration $t$, bound the $\ell_{2}$–distance between the clipped, averaged gradient updates under any admissible secret pair $(s_{i},s_{j})\in\mathcal{Q}$ and any coupling of their corresponding data-generating distributions.
@@
$\Pr[\mathcal{M}(X)\in E]\leq e^{\varepsilon_{g}}\Pr[\mathcal{M}(X^{\prime})\in E]+\delta_{g},\;\forall E\subseteq\mathcal{Y},\ \forall X\sim_{k}X^{\prime}.$ (43)
In DP-SGD, one typically fixes $k$ and uses a moments accountant to obtain $(\varepsilon_{g},\delta_{g})$ as a function of $k$, the sampling probability $q=B_{t}/N$, the number of steps $T$, and the per-step noise scale $\sigma_{\mathrm{DP}}$.
A natural question is whether group DP-SGD can be used as a proxy for SRPP-/sa-SRPP-SGD by choosing a group size $k$ that reflects the secret-induced discrepancy at iteration $t$. However, $k$ in group DP is a dataset-level Hamming distance (an integer), whereas our sa-HUC uses a minibatch-level mean-square discrepancy cap $\overline{K}_{t}^{2}$ satisfying $\mathbb{E}_{\eta_{t}}[\Delta_{t}^{2}]\leq\overline{K}_{t}^{2}$. A dimensionally compatible proxy would be $k\approx\sqrt{\overline{K}_{t}^{2}}$ (or more conservatively $k\approx\lceil\sqrt{\overline{K}_{t}^{2}}\rceil$), but even with such a choice, group DP-SGD does not generally yield a certified SRPP/sa-SRPP guarantee for an arbitrary Pufferfish scenario.
@@
  • There is no general result that, given a group-DP guarantee (43) for some group size $k$, produces an $(\alpha,\varepsilon)$-SRPP or sa-SRPP guarantee for an arbitrary Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$. In particular, even if one chooses a proxy group size $k\approx\lceil\sqrt{\overline{K}_{t}^{2}}\rceil$ (or uses a tail cap $k=K_{t}(\delta_{t})$) motivated by our HUC/sa-HUC analysis, there is no closed-form or universal mapping from $(\varepsilon_{g},\delta_{g},k)$ to an SRPP/sa-SRPP budget $\varepsilon$.
  • Even when we fix a specific Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$ and choose a proxy group size $k\approx\lceil\sqrt{\overline{K}_{t}^{2}}\rceil$ (or $k=K_{t}(\delta_{t})$), the resulting $(\varepsilon_{g},\delta_{g})$ group-DP-SGD guarantee is in general not a certified conservative bound for our $(\alpha,\varepsilon)$ SRPP/sa-SRPP guarantees, and there is no theorem ensuring that this group-DP guarantee upper-bounds the leakage measured in the SRPP/sa-SRPP sense.
G.1 No Universal Mapping
We first formalize the absence of a universal conversion from group-DP guarantees to SRPP/sa-SRPP guarantees.
@@ A DP-to-SRPP conversion map is a function
that, for given group-DP parameters $(\epsilon_{g},\delta_{g})$ and group size $k$, assigns an SRPP budget $F(\epsilon_{g},\delta_{g},k)$ with the following property: for every mechanism $\mathcal{M}$ and every Pufferfish scenario $(\mathcal{S},\mathcal{Q},\Theta)$, if $\mathcal{M}$ is $(\epsilon_{g},\delta_{g})$-group-DP with group size $k$, then for all $\theta\in\Theta$ and all $(s_{i},s_{j})\in\mathcal{Q}$ the SRPP/sa-SRPP divergence between the distributions of $\mathcal{M}$ under $(S=s_{i},\theta)$ and $(S=s_{j},\theta)$ is bounded by $F(\epsilon_{g},\delta_{g},k)$.
@@
Proposition G.1.
There is no function $F$ satisfying Definition G.1, even if we restrict to pure group-DP ($\delta_{g}=0$) and to finite secret sets $\mathcal{S}$.
Proof of Proposition G.1.
@@
Let the dataset space be $\mathcal{X}=\{X_{0},X_{1}\}$, and choose $k\in\mathbb{N}$ so that $d_{H}(X_{0},X_{1})>k$, where $d_{H}$ is the Hamming distance. Under the group adjacency relation restricted to $\mathcal{X}=\{X_{0},X_{1}\}$, there are no distinct adjacent pairs: the only pairs with $d_{H}(X,X^{\prime})\leq k$ are $(X_{0},X_{0})$ and $(X_{1},X_{1})$. Hence, for any mechanism $\mathcal{M}$ and any $(\epsilon_{g},\delta_{g})$, the group-DP inequality (43) is trivially satisfied, so $\mathcal{M}$ is $(\epsilon_{g},\delta_{g})$-group-DP for all choices of $(\epsilon_{g},\delta_{g})$ and $k$.
@@
In particular, the Rényi divergence of any order $\alpha>1$ between the distributions of $\mathcal{M}(X)$ given $S=0$ and $S=1$ is $\infty$ (i.e., the two output laws are mutually singular), so no finite SRPP budget can hold.
@@
Proposition G.1 remains true if, instead of allowing arbitrary mechanisms, we restrict attention to mechanisms obtained by running DP-SGD and calibrating noise for a fixed group size $k$ (including proxy choices such as $k=K_{t}(\delta_{t})$ or $k\approx\lceil\sqrt{\overline{K}_{t}^{2}}\rceil$ motivated by our HUC/sa-HUC analysis). We can always construct a Pufferfish scenario whose secrets are encoded in dataset differences at Hamming distance strictly greater than $k$ and hence outside the scope of the group-DP guarantee. Therefore, even when the group-DP parameters $(\epsilon_{g},\delta_{g},k)$ are known from a moments-accountant analysis of DP-SGD, there is no closed-form or universal transformation that converts them into the SRPP/sa-SRPP budget used in our HUC/sa-HUC–based SRPP-SGD calibration.
G.2 No Safe Bounds
We now specialize to the prevalence-based Pufferfish scenarios used in our SGD privatization experiments, where the secret changes only the label configuration on a fixed feature set (e.g., the cat label in CIFAR-10). Concretely, we consider two secrets $s_{0},s_{1}$ corresponding to two datasets $X^{(0)},X^{(1)}\in\bar{\mathcal{X}}^{N}$ with different prevalences $p_{\mathrm{low}}\neq p_{\mathrm{high}}$ of the protected class. Let
@@ G.2 No Safe Bounds
be the number of records whose labels differ between the two secrets. In our HUC/sa-HUC analysis, $\tilde{\Delta}$ influences the secret-induced update discrepancy through the (minibatch-level) discrepancy cap sequence, e.g., via $K_{t}$ (or $\mathbb{E}_{\eta_{t}}[K_{t}^{2}]$), which determines the SRPP/sa-SRPP noise calibration. A natural question is whether one can instead treat such quantities as an effective group size and calibrate standard group DP-SGD with some $k$ derived from $K_{t}$ or $\mathbb{E}_{\eta_{t}}[K_{t}^{2}]$. The next proposition shows that, even in prevalence-based scenarios, a group-DP guarantee with any fixed $k$ does not yield a general conservative (i.e., universally safe) bound on secret-level SRPP/sa-SRPP leakage.
Proposition G.2 (Group-DP does not upper-bound secret leakage in general).
Fix any group size $k\in\mathbb{N}$ and any group-DP parameters $(\varepsilon_{g},\delta_{g})$. There exist
@@
    a mechanism $\mathcal{M}$ that is $(\varepsilon_{g},\delta_{g})$-group-DP with group size $k$
such that, for every $\alpha>1$, the secret-level Rényi divergence (hence SRPP/sa-SRPP leakage) between the output laws under $(S=s_{0},\theta)$ and $(S=s_{1},\theta)$ can be made arbitrarily large. In particular, there is no finite function of $(\varepsilon_{g},\delta_{g},k)$ that universally upper-bounds this secret-level leakage.
@@
Fix $k\in\mathbb{N}$ and $(\varepsilon_{g},\delta_{g})$. Choose a dataset size $N>k$ and an integer $\tilde{\Delta}\in\{k+1,\dots,N\}$.
@@
Let the record domain be $\bar{\mathcal{X}}=\mathcal{A}\times\{0,1\}$, where $\mathcal{A}$ denotes non-sensitive features and the second coordinate is the protected-class label. Let the dataset space be $\mathcal{X}:=\bar{\mathcal{X}}^{N}$. Fix a feature vector collection $(a_{1},\dots,a_{N})\in\mathcal{A}^{N}$ and two label vectors $(b^{(0)}_{1},\dots,b^{(0)}_{N})$ and $(b^{(1)}_{1},\dots,b^{(1)}_{N})$ such that: (i) the empirical prevalences differ, i.e. $\frac{1}{N}\sum_{i=1}^{N}b^{(0)}_{i}=p_{\mathrm{low}}$ and $\frac{1}{N}\sum_{i=1}^{N}b^{(1)}_{i}=p_{\mathrm{high}}$ with $p_{\mathrm{low}}\neq p_{\mathrm{high}}$, and (ii) exactly $\tilde{\Delta}$ labels differ, i.e. $\{i:b^{(0)}_{i}\neq b^{(1)}_{i}\}=\tilde{\Delta}$. Define
$X^{(0)}:=\bigl((a_{i},b^{(0)}_{i})\bigr)_{i=1}^{N},\qquad X^{(1)}:=\bigl((a_{i},b^{(1)}_{i})\bigr)_{i=1}^{N},$
so that $d_{H}(X^{(0)},X^{(1)})=\tilde{\Delta}>k$.
Define the Pufferfish scenario by
$\mathcal{S}=\{s_{0},s_{1}\},\quad\Theta=\{\theta\},\quad\mathcal{Q}=\{(s_{0},s_{1}),(s_{1},s_{0})\},$
and let the conditional data law be deterministic:
$P_{\theta}(X=X^{(0)}\mid S=s_{0})=1,\qquad P_{\theta}(X=X^{(1)}\mid S=s_{1})=1.$
Group adjacency.
Let group adjacency be defined by Hamming distance at most $k$:
@@
Since $d_{H}(X^{(0)},X^{(1)})=\tilde{\Delta}>k$, the pair $(X^{(0)},X^{(1)})$ is not adjacent, and the group-DP constraint never compares $\mathcal{M}(X^{(0)})$ to $\mathcal{M}(X^{(1)})$.
Mechanism with group-DP but large secret leakage.
Fix any statistic $T:\mathcal{X}\to\mathbb{R}$ with bounded group sensitivity at size $k$,
$\Delta_{k}(T):=\sup_{X\sim_{k}X^{\prime}}|T(X)-T(X^{\prime})|<\infty,$
and choose $T$ so that $|T(X^{(0)})-T(X^{(1)})|=L$ for an arbitrary prescribed $L>0$ (e.g., define $T(X)=L\cdot\mathbf{1}\{X=X^{(0)}\}$; then $\Delta_{k}(T)\leq L$ because any $X^{\prime}\sim_{k}X^{(0)}$ must satisfy $X^{\prime}\neq X^{(1)}$). Let
$\mathcal{M}(X):=T(X)+Z,\qquad Z\sim\mathcal{N}(0,\sigma^{2}),$
with $\sigma$ calibrated so that $\mathcal{M}$ is $(\varepsilon_{g},\delta_{g})$-group-DP for group size $k$ under the adjacency $\sim_{k}$ (this is the standard Gaussian mechanism calibration based on $\Delta_{k}(T)$).
Under the above Pufferfish scenario, $\mathcal{M}(X)\mid(S=s_{0})$ is $\mathcal{N}(T(X^{(0)}),\sigma^{2})$ and $\mathcal{M}(X)\mid(S=s_{1})$ is $\mathcal{N}(T(X^{(1)}),\sigma^{2})$. For any $\alpha>1$, the Rényi divergence between these two output laws is
$\mathtt{D}_{\alpha}\!\bigl(\mathcal{N}(T(X^{(0)}),\sigma^{2})\,\|\,\mathcal{N}(T(X^{(1)}),\sigma^{2})\bigr)$
$=\frac{\alpha}{2(\alpha-1)}\cdot\frac{|T(X^{(0)})-T(X^{(1)})|^{2}}{\sigma^{2}}=\frac{\alpha}{2(\alpha-1)}\cdot\frac{L^{2}}{\sigma^{2}}.$
Since $L>0$ can be chosen arbitrarily large while the group-DP parameters $(\varepsilon_{g},\delta_{g},k)$ are fixed (by construction), this Rényi divergence (hence SRPP/sa-SRPP leakage) can be made arbitrarily large. Therefore no finite function of $(\varepsilon_{g},\delta_{g},k)$ can universally upper-bound secret-level leakage across such prevalence-based scenarios. ∎
@@
Proposition G.2 shows that even for prevalence-based secrets as in our CIFAR-10 experiments, calibrating DP-SGD using a group size chosen from HUC/sa-HUC-derived quantities (e.g., $k=\lceil\sqrt{\overline{K}_{t}^{2}}\rceil$ or $k=K_{\mathrm{cap}}$) does not yield a certified conservative bound in the SRPP/sa-SRPP sense. Group-DP is indexed by Hamming-adjacent dataset pairs, whereas SRPP/sa-SRPP quantify divergence between the secret-conditioned output laws in the specified Pufferfish scenario. Our HUC/sa-HUC-based SRPP-SGD calibration is therefore not a reparameterization of group DP-SGD, but a distinct, secret-aware analysis.
Appendix H Minimality and Attainability of HUC
In this section, we characterize the minimality and attainability of HUC defined by Definition 5.1 in Sec. 5.1.
Proposition H.1.